OpenStudy (anonymous):

if h(x) = x^3, show that h(a) = -h(-a)

Parth (parthkohli):

To do so, just calculate \(h(a)\) and \(-h(-a)\)...\[h(a) = a^3\]\[-h(-a)= -(-a)^3 = \cdots\]